Calculate the milliliters required to prepare 0.50 L of 0.1 N solution of HCl from a solution of HCl at 37.5% by weight
Answer #1

First, the molar concentration of the mother solution is calculated:

C1 = 37.5 g HCl / 100 g Sol * (1 g sol / 1 mL sol) * (1000 mL / 1 L) * (1 mol HCl / 36.46 g) = 10.29 M

The volume of solution 1 needed is calculated:

V1 = C2 * V2 / C1 = 0.1 M * 500 mL / 10.29 M = 4.9 mL

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.